# Neutrino oscillations originate from virtual excitation of Z bosons

### v2

@@ Abstract
To account for neutrino oscillations, it is postulated that the neutrino has nonvanishing mass and each flavor eigenstate is formed by three distinct mass eigenstates, whose probability amplitudes interfere with each other during its propagation. However, I find that the energy conservation law requires these mass eigenstates, if they exist, to be entangled with distinct joint energy eigenstates of the other particles produced by the same weak interaction as the neutrino. This entanglement destroys the quantum coherence among the neutrino’s mass eigenstates, which is responsible for flavor oscillations under the aforementioned postulation. I reveal that the neutrino oscillations actually originate from virtual excitation of the Z bosonic field diffusing over the space. During the propagation, the neutrino can continually excite and then immediately re-absorb a virtual Z boson. This virtual bosonic excitation produces a backaction on the neutrino, enabling it to oscillate among three flavors. When the neutrino propagates in matter, its behavior is determined by the competition between the coherent flavor transformation and decoherence effect resulting from scatterings.
@@
The discovery of neutrino flavor transformations represents one of the most striking advancements of physical science achieved during the past few decades [5-29]. Such phenomena are commonly believed to show that neutrinos have nonvanishing mass [8-31], which is beyond the standard model. To account for these phenomena, it is assumed that there exist three distinct neutrino mass eigenstates. Each flavor eigenstate corresponds to a quantum-mechanical superposition of these mass eigenstates, and vice versa [32]. During the propagation of the neutrino, the probability amplitude associated with each mass eigenstate accumulates a phase that is approximately proportional to the square of the corresponding mass. The time-evolving phase differences among these probability amplitudes give arise to flavor oscillations, as a consequence of their quantum interference. Such an interpretation, however, overlooks the crucial question: Does nature allow neutrinos produced by weak interactions to be in such superposition states? As all neutrinos that were observed to exhibit flavor transformations had been generated through weak interactions, only when the answer of this question is affirmative can the aforementioned interpretation be reasonable. However, I find the answer is negative for the following reason. If the neutrino (or antineutrino) produced by a weak interaction has three distinct mass eigenstates ($\left|\nu_{i}\right\rangle$), these eigenstates should be correlated with different joint states ($\left|\psi_{i}\right\rangle$) of the other particles (accompanying particles) produced by the same interaction, e.g., the joint states for the proton and electron produced by a $\beta$ decay. The energy differences of these joint states need to compensate for those of the neutrino mass eigenstates in order to conserve the total energy. As detailed below, this correlation would destroy quantum interference effects among the neutrino mass eigenstates.
Suppose that the entire system just before participating in the weak interaction (e.g., the neutron just before making a $\beta$ decay) has a definite total energy $E_{T}$. Then the entire state of the neutrino and accompanying particles produced by such an interaction can be written as $\mathop{\displaystyle\sum}\limits_{i}C_{i}\left|\nu_{i}\right\rangle\left|\psi_{i}\right\rangle$. The energy of $\left|\psi_{i}\right\rangle$ should be $E_{i}=E_{T}-\sqrt{p^{2}+m_{i}^{2}}$, where $p$ represents the momentum of the neutrino, and $m_{i}$ is the mass associated with $\left|\nu_{i}\right\rangle$. Only with such an arrangement can the final system has the same definite total energy, $E_{T}$, as the initial one, which ensures the process to comply with the energy conservation law. Under the assumption $m_{1}\neq m_{2}\neq m_{3}$, we have $E_{1}\neq E_{2}\neq E_{3}$, which implies that $\left|\psi_{1}\right\rangle$, $\left|\psi_{2}\right\rangle$, and $\left|\psi_{3}\right\rangle$ correspond to three energy eigenstates with different eigenvalues for the subsystem combined by the accompanying particles, thereby being orthogonal to each other. Due to this orthogonality, when the freedom degrees of the accompanying particles are traced out, the neutrino is left in a classically mixed state, described by the density operator, $\rho=\mathop{\displaystyle\sum}\limits_{i}\left|C_{i}\right|^{2}\left|\nu_{i}\right\rangle\left\langle\nu_{i}\right|$. Consequently, there does not exist any quantum coherence among the mass eigenstates, so that their mass differences cannot result in any flavor oscillation.
@@
As our theoretical model allows occurrence of oscillations even if the neutrino is massless, one may argue that the state of a massless particle cannot be changed during its propagation in the vacuum. This is the case if the particle remains unperturbed, as exemplified by the free propagation of a photon. However, the photonic state can be changed by some scattering process. For example, a photon can transform between modes $a$ and $b$ with the assistance of $\Lambda$-type three-level atoms, which have an excited state $\left|e\right\rangle$ and two ground states $\left|g\right\rangle$ and $\left|f\right\rangle$ [38]. The transitions $\left|e\right\rangle\rightarrow\left|g\right\rangle$ and $\left|e\right\rangle\rightarrow\left|f\right\rangle$ are respectively coupled to these two modes with the same detuning. When the detuning is much larger than the coupling strengths, the photon would oscillate between these two modes through a Raman scattering process, mediated by virtual excitation of the atoms. The flavor oscillations of the neutrino bear some similarity to such a Raman process. As the neutrino is inevitably disturbed by the vacuum fluctuations of the gauge fields pervading the space, its propagation cannot be really free even in the vacuum. The backaction of the virtually excited Z gauge field on the neutrino is analogous to that of the virtually excited atoms on the photon in the Raman process. One may further ask: Does there exist any other mechanism that can mediate the neutrino flavor transformation? The answer is presumably negative for the following reason. The W and Z bosonic fields are the only elements that can interact with the neutrino during its propagation in the vacuum. As the neutrino flavor is defined by the charged current that is mediated by the W bosonic field, the vacuum fluctuation of the Z bosonic field seems to be the only mechanism that can change the neutrino flavor.
This work was supported by the National Natural Science Foundation of China under Grant No. 12274080.

### v3

@@ Abstract
To account for neutrino oscillations, it was postulated that the neutrino has nonvanishing mass and each flavor eigenstate is formed by three distinct mass eigenstates, whose probability amplitudes interfere with each other during its propagation. However, for a neutrino produced by the decay of an unstable particle, these mass eigenstates, if they exist, would be entangled with distinct joint momentum eigenstates of the entire system, composed of all the particles produced by the decay. This entanglement destroys the quantum coherence among the neutrino’s mass eigenstates, which is responsible for flavor oscillations under the aforementioned postulation. I reveal that the neutrino oscillations actually originate from virtual excitation of the Z bosonic field diffusing over the space. During the propagation, the neutrino can continually excite and then immediately re-absorb a virtual Z boson. This virtual bosonic excitation produces a backaction on the neutrino, enabling it to oscillate among three flavors. When the neutrino propagates in matter, its behavior is determined by the competition between the coherent flavor transformation and decoherence effect resulting from scatterings.
@@
The discovery of neutrino flavor transformations represents one of the most striking advancements of physical science achieved during the past few decades [5-29]. Such phenomena are commonly believed to show that neutrinos have nonvanishing mass [8-31], which is beyond the standard model. To account for these phenomena, it is assumed that there exist three distinct neutrino mass eigenstates. Each flavor eigenstate corresponds to a quantum-mechanical superposition of these mass eigenstates, and vice versa [32]. During the propagation of the neutrino, the probability amplitude associated with each mass eigenstate accumulates a phase that is approximately proportional to the square of the corresponding mass. The time-evolving phase differences among these probability amplitudes give arise to flavor oscillations, as a consequence of their quantum interference. However, as has been proved [33,34], if the state of a neutrino produced from the decay of an unstable particle consists of different mass eigenstates, the mass degree of freedom of the neutrino is necessarily entangled with the momentum freedom degrees of the accompanying particles or/and that of the neutrino itself, irrespective of the momentum distribution of the unstable particle. Due to this entanglement, the neutrino is essentially in a classical mixture of these mass eigenstates when all the momentum freedom degrees are traced out. Consequently, the neutrino cannot has a definite flavor at the production, and the population of each flavor eigenstate cannot be changed during the propagation.
@@
As our theoretical model allows occurrence of oscillations even if the neutrino is massless, one may argue that the state of a massless particle cannot be changed during its propagation in the vacuum. This is the case if the particle remains unperturbed, as exemplified by the free propagation of a photon. However, the photonic state can be changed by some scattering process. For example, a photon can transform between modes $a$ and $b$ with the assistance of $\Lambda$-type three-level atoms, which have an excited state $\left|e\right\rangle$ and two ground states $\left|g\right\rangle$ and $\left|f\right\rangle$ [40]. The transitions $\left|e\right\rangle\rightarrow\left|g\right\rangle$ and $\left|e\right\rangle\rightarrow\left|f\right\rangle$ are respectively coupled to these two modes with the same detuning. When the detuning is much larger than the coupling strengths, the photon would oscillate between these two modes through a Raman scattering process, mediated by virtual excitation of the atoms. The flavor oscillations of the neutrino bear some similarity to such a Raman process. As the neutrino is inevitably disturbed by the vacuum fluctuations of the gauge fields pervading the space, its propagation cannot be really free even in the vacuum. The backaction of the virtually excited Z gauge field on the neutrino is analogous to that of the virtually excited atoms on the photon in the Raman process. One may further ask: Does there exist any other mechanism that can mediate the neutrino flavor transformation? The answer is presumably negative for the following reason. The W and Z bosonic fields are the only elements that can interact with the neutrino during its propagation in the vacuum. As the neutrino flavor is defined by the charged current that is mediated by the W bosonic field, the vacuum fluctuation of the Z bosonic field seems to be the only mechanism that can change the neutrino flavor. Finally, it should be noted that the virtual excitation of Z bosons cannot lead to flavor transformations between different flavors of charged leptons owing to their large mass differences.